# Eleanor Wiske Dillon Sonia Jaffe Microsoft Research Microsoft Research Nicole Immorlica Christopher T. Stanton Microsoft Research Harvard Business School, NBER, and CEPR

### v1

@@ 2.1 Methods
where $\delta_{m_{t}}$ captures calendar year-month effects, $\tau_{ft}$ is the event month for firm $f$ (month relative to when the firm started the experiment), and $Z_{i}$ is the worker’s experimental assignment. We include both calendar time and event time controls because firms entered and exited the study on a rolling basis. Note that while outcomes are measured on a weekly basis, we capture time-varying effects at a monthly frequency to reduce noise. We use Newey-West standard errors with a lag of 9 weeks to handle autocorrelation in outcomes.
@@
As highlighted in the previous section, many workers who had access to Copilot did not use it regularly, and some workers who were assigned to the control group nonetheless obtained Copilot licenses. To adjust for low usage and non-compliance we also follow Duflo (2001) and Hudson, Hull and Liebersohn (2017) and estimate an instrumented difference-in-difference specification. In this case, we use treatment assignment as an instrument for whether a worker was an above-median Copilot user, using Copilot in at least 42% of the post-treatment weeks. The single treatment effect regression equations are [^11] Note that $D_{it}$ does not vary (for an individual) across weeks within the pre- or post- period. It is one for all post-period weeks for individuals who were above-median Copilot users and zero otherwise. Because of the time fixed effects, $\hat{D}_{it}$ can vary month-to-month.
$Y_{it}$ $=\alpha_{i}+\delta_{m_{t}}+\gamma_{f,\tau_{ft}}+\beta\hat{D}_{it}+\epsilon_{it}$ (2)
$\hat{D}_{it}$ $=\tilde{\alpha}_{i}+\tilde{\delta}_{m_{t}}+\tilde{\gamma}_{f,\tau_{ft}}+\tilde{\beta}Z_{i}\cdot\mathbf{1}\{\tau_{ft}\geq 0\}+\eta_{it}.$
These estimates capture the local average treatment effect of Copilot for the complying workers who actually use it. To extend this IV setup to the multi-period event studies, we instrument each $\hat{D}_{it}$ with interactions of $Z_{i}$ and event time.
@@ 2.2 Average Effects
During this early adoption phase, the most prominent impact of Copilot on worker behavior was changes in how workers managed their email. In the pre-period, the average worker in our sample spent over 11 hours per week reading and responding to emails [^12] Details on the construction of all worker behavior measures are in Appendix B.1. , more than a quarter of a typical work week. For most workers, reading and writing emails does not represent their core job responsibilities but rather a communication overhead cost that supports their productive work. On average, workers with access to Copilot spent 1.4 hours less per week working on email in Outlook and workers who used Copilot regularly spent almost 3 fewer hours per week, a 25% decline from their pre-period average. Figure 2 illustrates that this reduction in Outlook time is apparent even in the first month after workers got access to the tool and persists throughout the study.
@@
Note: These plots show the coefficients from regressions similar to Equations (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates, where the $\beta$’s are allowed to vary by firm’s month from Copilot adoption. Vertical lines show 95% confidence intervals based on Newey-West standard errors with a lag of 9 weeks. Horizontal dashed lines give the average estimated effect from Table 2.
@@
Table 2: Effect of Copilot
@@
Outlook total session time 11.53 -1.35** -2.95** 5,803 281,283
@@
Total concentration time 24.61 0.90** 2.05** 5,620 247,907
@@
Outlook out-of-hours session time 2.08 -0.30** -0.65** 5,803 281,283
@@
Unique email threads replied to 13.22 -0.23* -0.47* 6,116 325,472
@@
Time to reply from email delivery 14.77 -0.31** -0.61** 6,111 285,332
@@
Teams meeting time 5.21 0.041 0.081 5,698 208,754
@@
Share of meetings attended late/early 0.30 0.006** 0.012** 5,645 178,218
@@
Recurring Teams meetings time 2.32 0.013 0.025 5,698 254,319
@@
Word total session time 1.35 0.018 0.032 2,150 114,863
@@
Avg. document time to complete 175.56 -6.93 -11.65 2,150 36,523
@@
Completed collaborative documents 0.15 0.011 0.019 2,150 91,213
@@
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Column 1), the intent-to-treat coefficients for the effect of Copilot (Column 2) and the LATE coefficients (Column 3). Cross-user standard deviations are estimates of the standard deviation of the random intercept in the model $Y_{it}=\mu_{i}+\epsilon_{it}$. All time metrics are in hours. Share of meetings attended late/early refers to the share of Teams meetings that a worker joined more than 5 minutes after the meeting started or left more than 5 minutes before it ended. We assign documents to a worker if they were the ‘primary editor’, making the more most edits to the document.
@@
Looking in more detail at how workers changed their behavior when using Copilot, we find more evidence for the second hypothesis: workers were more likely to adjust behaviors that could be changed unilaterally and less likely to shift behaviors that required coordination with colleagues. In the top panel of Table 2, we see that workers’ responsiveness and intensity of email communication with colleagues is roughly unchanged. Workers who used Copilot replied to roughly the same number of conversations per week (13 vs. 13.5) and replied to emails slightly faster than before. These workers simply take less time to do roughly the same quantity of email work. To illustrate some of the benefits of this shift in behavior, note that workers who use Copilot not only spending less time in Outlook each week but also condense their work into fewer sessions. This opens up two more hours per week of concentration time, blocks of 30 minutes or more during the regular workday with no Outlook activity. Workers using Copilot also cut back on time spent in email outside of their regular work hours.
@@
If the null effect on total time spent in meetings was masked by offsetting shifts in how workers were using meetings, we would expect to see some changes in the types of meetings workers attended after gaining access to Copilot. However, as shown in Table 2, we find no impact on time spent in recurring meetings relative to meetings overall. [^13] As shown in appendix table A3, we also find no significant effect on the number of meetings attended, ruling out a shift toward more shorter meetings, and on most meetings by type (long meetings with few participants, short meetings with many participants, etc.), which previous research has shown correlate with meeting purpose. While we do estimate small but statistically significant overall increases in large meetings for workers with Copilot access, the event studies in Figure A5 show that this increase is concentrated in the first month of Copilot access, which suggests these might be required training sessions in the new tool. The only impact of Copilot on meeting-related behavior that we find is a small, precisely estimated increase in the share of meetings where treated workers join late or leave early, an individual choice. Teams is the most common application in which workers used Copilot, so they likely found value in using generative AI in this context, perhaps in lowering effort to retain information from meetings or providing the ability to focus on participating instead of taking notes. However, this value did not translate into any visible shift in the structure of work during our sample.
@@
Focusing on workers who were regular users of Word in the pre-period, the bottom panel of Table 2 provides more details on how Copilot access and use affects workers’ document production. We do not have the power to measure these effects very precisely because over half of study participants were not regular Word users. However, there are some suggestive patterns. Consistent with Noy and Zhang (2023), we find that workers with Copilot complete their documents more quickly, although the magnitude of the effect is far smaller than what they found in the lab, and we cannot reject the null hypothesis of no effect. Interestingly, the time savings are strongest, in both time saved (over 2 days) and percent reductions (20%), in documents where our studied worker was the primary editor but at least one other coworker also contributed. These multi-author documents take longer to complete, nearly 12 days on average, so there may be more scope for time savings. Workers may also feel more comfortable using generative AI staying on top of colleagues’ edits and outstanding tasks than drafting text in work documents. The larger effects in collaborative documents also reassure us that we are seeing some behavior shift beyond casual experimentation with a new tool. Consistent with workers changing the way they do their own work, but not taking on different responsibilities, we find no change in the number of documents where the studied workers are the primary editor.
2.3 Effects by Peer Copilot Use
@@
Outlook total session time 11.77 (5.98) 11.30 (6.78) -1.50** (0.095) -1.27** (0.10) -3.04** (0.19) -2.97** (0.23) 140,463 140,820
Total concentration time 24.53 (5.86) 24.69 (6.57) 1.05** (0.084) 0.78** (0.087) 2.22** (0.18) 1.90** (0.21) 128,272 119,635
Outlook out-of-hours session time 2.12 (1.55) 2.04 (1.75) -0.33** (0.026) -0.28** (0.027) -0.67** (0.053) -0.64** (0.062) 140,463 140,820
Unique email threads replied to 13.60 (11.43) 12.86 (11.38) -0.26 (0.11) -0.26 (0.12) -0.49 (0.21) -0.58 (0.26) 160,538 164,934
Time to reply from email delivery 15.00 (4.82) 14.54 (4.81) -0.25 (0.13) -0.35* (0.13) -0.47 (0.24) -0.75* (0.29) 144,189 141,143
@@
Teams meeting time 5.71 (3.13) 4.72 (3.34) 0.009 (0.047) 0.067 (0.048) 0.017 (0.087) 0.14 (0.10) 105,376 103,378
% Teams meetings attended late/early 0.28 (0.096) 0.31 (0.12) 0.006 (0.003) 0.006 (0.003) 0.010 (0.004) 0.013 (0.006) 92,260 85,958
Recurring Teams meetings time 2.51 (1.62) 2.13 (1.74) -0.008 (0.022) 0.031 (0.022) -0.014 (0.041) 0.066 (0.046) 130,249 124,070
@@
Word total session time 1.31 (0.91) 1.38 (0.95) 0.010 (0.031) 0.011 (0.033) 0.017 (0.055) 0.020 (0.061) 59,662 55,201
Avg. document time to complete 179.17 (86.74) 171.75 (77.25) -1.95 (7.69) -13.26 (7.71) -3.23 (12.74) -22.97 (13.35) 19,382 17,141
Avg. collaborative document TTC 280.76 (121.48) 275.15 (104.04) -36.29 (19.21) -24.95 (22.51) -54.44 (28.83) -42.13 (37.96) 6,507 4,526
Avg. non-collaborative document TTC 145.01 (77.50) 142.38 (77.72) 0.97 (7.48) -9.62 (7.46) 1.64 (12.63) -16.58 (12.85) 16,138 14,837
Completed documents 0.78 (0.75) 0.77 (0.75) -0.005 (0.023) 0.057 (0.025) -0.009 (0.041) 0.11 (0.046) 47,551 43,662
Completed collaborative documents 0.17 (0.23) 0.13 (0.21) 0.006 (0.008) 0.014 (0.007) 0.011 (0.014) 0.027 (0.014) 47,551 43,662
@@
Note: This table replicates the analysis in Table 2 separately for workers who had above or below median share of coworkers with Copilot licenses in the post-period. See that table not for explanations and variable definitions.
Table 3 splits our sample by whether our studied workers have above or below median share (12.5%) of close coworkers with access to Copilot. We see almost no meaningful differences in behavior between these two groups, with one exception in document creation. Workers who have fewer colleagues with access to Copilot author 14% more documents when they use Copilot themselves. These documents are primarily solo authored, since we see no increase in collaborative documents, and these solo documents are completed more quickly. This pattern suggests that some teams adapted in the short run and reallocated writing work to the one team member with access to the new generative AI tool. However, we do not expect this pattern to persist in the long run as generative AI tools become more widespread. Appendix Table A5 shows a similar exercise, splitting our sample into workers at firms with higher or lower overall Copilot adoption. Again, while there are some small differences in impacts across these two groups, we do not see evidence that firms with higher Copilot adoption showed more movement in behaviors that require group coordination. Overall, the lack of large differences suggests that the barriers to broader shifts in work patterns are not simply coordination costs within small teams but might include factors such as the time required to reimagine work flows and the institutional effort to drive these changes.
@@ 3 Conclusion
This study provides some of the earliest signals of how this individual use of generative AI is beginning to shape the patterns of work for knowledge workers. We observed workers during the first year in which generative AI tools were beginning to gain broad use at work. While their colleagues may have been using other generative AI tools, the workers we studied were often the only members of their team with access to Copilot during the early roll-out period. The changing behaviors we see are consistent with workers independently exploring these new tools and saving time on individual tasks, but the individual rollout did not exhibit the organization-wide shifts we might expect as teams and organizations begin larger transformations in response to this technology. We therefore expect that the patterns we find are only the beginning of a new era of work.

### v2

@@ 2.1 Methods
where $\delta_{m_{t}}$ captures calendar year-month effects, $\tau_{ft}$ is the event month for firm $f$ (month relative to when the firm started the experiment), and $Z_{i}$ is the worker’s experimental assignment. We include both calendar time and event time controls because firms entered and exited the study on a rolling basis. Note that while outcomes are measured on a weekly basis, we capture time-varying effects at a monthly frequency to reduce noise. We cluster standard errors at the firm level.
@@
As highlighted in the previous section, over 90% of workers who had access to Copilot used it at least once. The ITT estimates are thus very close to an IV estimate where any use of Copilot over 6 months is the measure of takeup or adoption. However, many treated workers did not use Copilot regularly. In addition, some workers who were assigned to the control group nonetheless obtained Copilot licenses. To understand patterns of adoption that reflect more habitual and intense usage, we present an instrumented difference-in-difference specification (Duflo, 2001; Hudson, Hull and Liebersohn, 2017). In this case, we use treatment assignment as an instrument for whether a worker used Copilot in at least 50% of the post-treatment weeks, which allows us to understand effects for compliers who are induced to use Copilots regularly. [^11] Note that $D_{it}$ does not vary (for an individual) across weeks within either the pre- or post- period. It is one for all post-period weeks for individuals who were above-50% Copilot users and zero otherwise. The outcome equation and first stage are, respectively:
$Y_{it}$ $=\alpha_{i}+\delta_{m_{t}}+\gamma_{f,\tau_{ft}}+\beta D_{it}+\epsilon_{it}$ (2)
$D_{it}$ $=\tilde{\alpha}_{i}+\tilde{\delta}_{m_{t}}+\tilde{\gamma}_{f,\tau_{ft}}+\tilde{\beta}Z_{i}\cdot\mathbf{1}\{\tau_{ft}\geq 0\}+\eta_{it}.$
These estimates capture the local average treatment effect of Copilot for the complying workers who actually use it. To extend this IV setup to the multi-period event studies, we interact $D_{it}$ with event month and instrument with interactions of $Z_{i}$ and event month.
Because we test multiple hypotheses, we employ sharpened $q$-values that account for the increased False Discovery Rate as the number of tests increases (Anderson, 2008). The $q$-values (reported as asterisks in tables) can be interpreted similarly to a $p$-value after implementing a correction that accounts for the total number of hypotheses tested in the analysis.
@@ 2.2 Average Effects
We are able to measure two broad sets of outcome measures: time use in digital applications and quantity-related metrics, like the number of meetings attended, documents created, etc. The quantity measures are straightforward. The time use measures for Outlook and Word are constructed from timestamps associated with user actions. Specifically, we measure time use in Outlook and Word using the concept of a session, which captures a block of time associated with an application’s use. For example, an Outlook session begins when a user opens an email in the reading pane. The session ends when either: a) the user exits the reading pane and does not open another email for 15 minutes, or b) 15 minutes elapses and no other email has been opened. This stopping rule is applied to subsequent emails after a session begins. A similar concept is applied to calculate session time in Word, where sessions are based on 15 minute gaps in activities (e.g., interactions with a document or series of documents). [^12] Details on the construction of all worker behavior measures are in Appendix B.1.
During this early adoption phase, the most prominent impact of Copilot on worker behavior involved changes in how workers managed their email. In the pre-period, the average worker in our sample spent 11.7 hours per week reading and responding to emails, more than a quarter of a typical work week. For most workers, reading and writing emails does not represent their core job responsibilities but rather a communication overhead cost that supports their productive work. On average, workers with access to Copilot spent 1.3 hours less per week working on email in Outlook and workers who used Copilot regularly spent 3.6 fewer hours per week, a 31% decline from their pre-period average. Figure 2 illustrates that this reduction in Outlook time is apparent even in the first month after workers got access to the tool and persists throughout the study.
@@
Note: These plots show the coefficients from regressions similar to Equations (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates, where the $\beta$’s are allowed to vary by firm’s month from Copilot adoption. Vertical lines show 95% confidence intervals based on firm-clustered errors. Horizontal dashed lines give the average estimated effect from Table 2.
@@
Table 2: Effects of Copilot Provision and Takeup
@@
Outlook total session time 11.65 -1.29** -3.56** 6,441 313,461
@@
Total concentration time 26.18 1.44** 3.96** 6,441 313,461
@@
Outlook out-of-hours session time 2.19 -0.30** -0.83** 6,441 313,461
@@
Unique email threads replied to 13.56 -0.095 -0.24 7,137 375,642
@@
Time to reply from email delivery 16.43 -0.19 -0.46 7,095 320,811
@@
Teams meeting time 5.22 0.081 0.19 6,170 221,116
@@
Share of meetings attended late/early 0.31 0.005* 0.011 6,101 186,486
@@
Recurring Teams meetings time 2.32 0.025 0.058 6,170 268,353
@@
Word total session time 1.63 0.11 0.24 2,525 132,213
@@
Avg. document time to complete 186.54 -4.93 -9.80 2,525 40,366
@@
Completed collaborative documents 0.14 0.012* 0.026 2,525 104,438
@@
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Column 1), the intent-to-treat coefficients for the effect of Copilot (Column 2) and the LATE coefficients for more frequent Copilot use (Column 3). Cross-user standard deviations are estimates of the standard deviation of the random intercept in the model $Y_{it}=\mu_{i}+\epsilon_{it}$. All time metrics are in hours. Share of meetings attended late/early refers to the share of Teams meetings that a worker joined more than 5 minutes after the meeting started or left more than 5 minutes before it ended. We assign documents to a worker if they were the ‘primary editor’, making the most edits to the document. Standard errors are clustered by firm. Asterisks reference sharpened $q$-values that adjust for the false discovery rate, as in Anderson (2008).
@@
Looking in more detail at how workers changed their behavior when using Copilot, we find more evidence for the second hypothesis: workers were more likely to adjust behaviors that could be changed unilaterally and less likely to shift behaviors that required coordination with colleagues. In the top panel of Table 2, we see that workers’ responsiveness and intensity of email communication with colleagues is roughly unchanged. Workers who used Copilot regularly replied to roughly the same number of conversations per week (13.8 vs. 14.1) and, if anything, replied to emails slightly faster, though the difference is not significant. These workers simply take less time to do roughly the same quantity of email work. To illustrate some of the benefits of this shift in behavior, note that workers who use Copilot not only spend less time in Outlook each week but also condense their work into fewer sessions. This opens up almost 4 hours per week of concentration time, blocks of 30 minutes or more during the regular workday with no Outlook activity. Workers using Copilot also cut back on time spent in email outside of their regular work hours.
@@
If the null effect on total time spent in meetings was masking offsetting shifts in how workers were using meetings, we would expect to see changes in the types of meetings workers attended after gaining access to Copilot. However, as shown in Table 2, we find no impact on time spent in recurring meetings. [^13] As shown in appendix table A4, we find statistically significant but small (5%) increases in the count of recurring meetings and few changes in meeting type (long meetings with few participants, short meetings with many participants, etc.), which previous research has shown correlate with meeting purpose. As shown in the event studies in Figure A5, the increased participation in large meetings is concentrated in the first month of Copilot access, which suggests these might be required training sessions in the new tool. Workers with access to Copilot are slightly more likely to join meetings late or leave early, an individual choice (the IV analysis loses statistical signficance). Teams is the most common application in which workers used Copilot, so they likely found value in using generative AI in this context, perhaps in lowering effort to retain information from meetings or providing the ability to focus on participating instead of taking notes. However, this value did not translate into any visible shift in the structure of work during our sample.
@@
Focusing on workers who were regular users of Word in the pre-period, the bottom panel of Table 2 provides more details on how Copilot access and use affects workers’ document production. We do not have the power to measure these effects very precisely because over half of study participants were not regular Word users. However, there are some suggestive patterns. Consistent with Noy and Zhang (2023), we find that workers with Copilot complete their documents more quickly, although the magnitude of the effect is far smaller than what they found in the lab, and we cannot reject the null hypothesis of no effect. Interestingly, the time savings are strongest, in both time to complete (over 3 days) and percent reductions (25%), in documents where our studied worker was the primary editor but at least one other coworker also contributed. These multi-author documents take longer to complete, 12 days on average, so there may be more scope for time savings. Workers may also feel more comfortable using generative AI to stay on top of colleagues’ edits and outstanding tasks compared to drafting text in work documents. The larger effects in collaborative documents also reassure us that we are seeing some behavior shift beyond casual experimentation with a new tool. Consistent with workers changing the way they do their own work, but not taking on different responsibilities, we find no change in the number of documents completed where the studied workers are the primary editor.
2.3 Effects by Peer Copilot Access
@@
Outlook total session time 11.84 (6.37) 11.45 (7.30) -1.43** (0.25) -1.16** (0.21) -3.69** (0.57) -3.42** (0.57) 168,591 144,870
Total concentration time 25.83 (6.21) 26.59 (7.34) 1.66** (0.27) 1.23** (0.22) 4.27** (0.64) 3.62** (0.64) 168,591 144,870
Outlook out-of-hours session time 2.21 (1.75) 2.18 (1.94) -0.32** (0.067) -0.28** (0.071) -0.82** (0.15) -0.84** (0.17) 168,591 144,870
Unique email threads replied to 14.54 (13.93) 12.58 (13.68) -0.21 (0.18) -0.047 (0.21) -0.51 (0.42) -0.13 (0.57) 189,426 186,216
Time to reply from email delivery 16.67 (5.30) 16.18 (5.38) -0.14 (0.15) -0.23 (0.20) -0.32 (0.34) -0.60 (0.53) 170,030 150,781
@@
Teams meeting time 5.79 (3.52) 4.63 (3.71) 0.029 (0.069) 0.13 (0.055) 0.063 (0.15) 0.32 (0.13) 116,333 104,783
% Teams meetings attended late/early 0.29 (0.11) 0.33 (0.13) 0.005 (0.003) 0.006 (0.004) 0.010 (0.006) 0.013 (0.010) 101,607 84,879
Recurring Teams meetings time 2.51 (1.86) 2.11 (2.01) -0.018 (0.030) 0.067 (0.032) -0.039 (0.065) 0.16 (0.072) 143,514 124,839
@@
Word total session time 1.61 (1.55) 1.65 (1.62) 0.12 (0.082) 0.088 (0.14) 0.25 (0.17) 0.20 (0.32) 72,324 59,889
Avg. document time to complete (TTC) 191.45 (98.98) 180.79 (86.00) -5.37 (8.45) -5.61 (10.83) -10.57 (16.73) -11.44 (21.83) 22,752 17,614
Avg. collaborative document TTC 294.49 (136.75) 277.91 (106.21) -54.44* (16.87) -18.22 (25.15) -94.40* (30.75) -32.43 (43.57) 7,399 4,437
Avg. non-collaborative document TTC 158.62 (85.60) 157.43 (94.14) 1.94 (10.76) -1.57 (11.07) 3.91 (21.63) -3.23 (22.76) 19,026 15,357
Completed documents 0.77 (0.81) 0.75 (0.80) 0.019 (0.032) 0.039 (0.036) 0.040 (0.066) 0.086 (0.080) 57,419 47,019
Completed collaborative documents 0.15 (0.22) 0.11 (0.20) 0.011 (0.008) 0.013 (0.007) 0.023 (0.016) 0.028 (0.015) 57,419 47,019
@@
Note: This table replicates the analysis in Table 2 separately for workers who had above- or below-median shares of coworkers with Copilot licenses in the post-period. See that table note for explanations and variable definitions. Standard errors are clustered by firm and $q$-values are calculated using adjustments for all hypotheses tested. The $q$-values for testing that the coefficients between the two groups are different all show non-statistically significant differences.
Table 3 splits our sample by whether our studied workers have above or below median share (11.9%) of close coworkers with access to Copilot. [^14] Given the average worker has 7 coworkers and any worker has at most 15 coworkers, many above-median workers have just one coworker with Copilot, the median is 2. We hypothesize that more intense treatment at the team level could spur larger shifts in coordinated work behavior, for example changing meeting behavior or reallocating primary editor responsibilities across workers. However, we find no such patterns in this sample. For collaborative documents, the decrease in time to complete is larger (and statistically signficant) for the above-median group, but the difference relative to the below-median group is not statistically significant. This lack of significant differences could suggest that firm strategies, rather than local team coordination, is the primary driver of changing practices in response to new technology. However, we note that even our “high saturation” teams have, on average, only 37% of coworkers with access to Copilot, so we may simply be unable to detect team-level effects from this experimental design.
@@ 3 Conclusion
This study provides some of the earliest signals of how this individual use of generative AI is beginning to shape the patterns of work for knowledge workers. We observed workers during the first year in which generative AI tools were beginning to gain broad use in firms. While their colleagues may have been using other generative AI tools, the workers we studied were often the only members of their team with access to Copilot during the early roll-out period. The changing behaviors we see are consistent with workers independently exploring these new tools and saving time on individual tasks, but the individual rollout did not exhibit the organization-wide shifts we might expect as teams and organizations begin larger transformations in response to this technology. We therefore expect that the patterns we find are only the beginning of a new era of work.